# Ontology Neural Network and ORTSF: A Framework for Topological Reasoning and Delay-Robust Control

### v1

@@ VII-B Delay-Compensated Control: Analytical Bound and Comparative Analysis
For delay system dynamics, the phase margin satisfies:
$\phi_{\operatorname{margin}}^{\operatorname{effective}}=\phi_{\operatorname{design}}-360f_{c}\Delta t+\phi_{\operatorname{comp}}$ (69)
where:
@@
In an ideal design:
$\phi_{\operatorname{comp}}\approx 360f_{c}\Delta t$ (71)
but practical compensators provide partial phase recovery.
@@
  • ORTSF: Predictive reasoning dynamically compensates delay, maintaining margin within safe thresholds ($\phi_{\operatorname{safe}}=20^{\circ}$).
@@ VIII-B Empirical Validation and Quantitative Insights
  • Figure 1: PH distance decay (final value $<0.05$)
@@ VIII-E Future Research Directions
  • Algorithmic acceleration: Aim for $<5$ ms/frame latency via approximations + GPU.
@@ A-A SLAM Optimization Cost Function
×$\quad\times\exp\Bigg{(}-\frac{1}{2}(z_{i}-h(x_{i},m_{i}))^{T}$
$\qquad\Sigma_{i}^{-1}(z_{i}-h(x_{i},m_{i}))\Bigg{)}$
@@
$=\prod_{i}\frac{1}{\sqrt{(2\pi)^{d}|\Sigma_{i}|}}$
×$\quad\times\exp\left(-\frac{1}{2}(z_{i}-h(x_{i},m_{i}))^{T}\Sigma_{i}^{-1}(z_{i}-h(x_{i},m_{i}))\right)$
@@ A-D Ontology
where $C$=classes, $P$=properties, $R$=relations.

### v2

@@ VII-B Delay-Compensated Control: Analytical Bound and Comparative Analysis
From our robust phase margin analysis (Patch-D), accounting for frequency drift and uncertainties:
$\phi_{\operatorname{margin}}^{\operatorname{effective}}$ $=\phi_{\operatorname{design}}-360(f_{c}+\Delta f_{c})\Delta t+\phi_{\operatorname{comp}}-\epsilon$ (92)
$\geq\phi_{\operatorname{safe}}+\sigma$ (93)
where $|\Delta f_{c}|\leq\alpha\|\Delta G\|$ accounts for crossover frequency drift, $\epsilon\geq 0$ represents modeling uncertainties, and $\sigma>0$ provides safety buffer. The compensator phase advance:
@@
must overcome both nominal delay and uncertainty effects.
@@
  • ORTSF: Discrete finite-difference predictor (Patch-B) with robust compensator design maintains margin within conservative safety bounds ($\phi_{\operatorname{safe}}+\sigma=20\textdegree+10\textdegree=30\textdegree$).
@@ VIII-B Empirical Validation and Quantitative Insights
  • Figure 1: PH distance decay (final value <0.05<0.05)
@@ VIII-E Future Research Directions
  • Algorithmic acceleration: Aim for <5<5 ms/frame latency via approximations + GPU.
@@ A-A SLAM Optimization Cost Function
$\quad\times\exp\Bigg{(}-\frac{1}{2}(z_{i}-h(x_{i},m_{i}))^{T}$ (119)
$\quad\quad\Sigma_{i}^{-1}(z_{i}-h(x_{i},m_{i}))\Bigg{)}$ (120)
@@
$=\prod_{i}\frac{1}{\sqrt{(2\pi)^{d}|\Sigma_{i}|}}$ (122)
$\quad\times\exp\Big{(}-\frac{1}{2}(z_{i}-h(x_{i},m_{i}))^{T}$ (123)
$\quad\quad\Sigma_{i}^{-1}(z_{i}-h(x_{i},m_{i}))\Big{)}$ (124)
@@ A-D Ontology
where CC=classes, PP=properties, RR=relations.